OpenStudy (anonymous):

How do I multiply this expression? 5(w + 5)/2w + 16 * 8(w + 8)/w - 12

Nnesha (nnesha):

\[\huge\rm \frac{ 5(w+5) }{ \color{green}{2w+16 }} \times \frac{ 8(w+8) }{ w-12 }\] first factor (2w+16)